OpenStudy (anonymous):

if sin (b) is sqareroot (2) / 2 what is cos (b) and tan (b) and cot(b)?

OpenStudy (anonymous):

cos sqrt 2 /2, tan 1 cot 1

OpenStudy (anonymous):

can u help with more?

OpenStudy (anonymous):

if tan (b) = squareroote (5)/ 2 what is cot(b) what is sin (b) and what is cos (b)

OpenStudy (anonymous):

and also if csc(V) = (4*squareroote 7)/ 7 then what is sin (V) and cos (V) and tan (V)

OpenStudy (anonymous):

so sorry for this. my teacher was absent for 2 weeks and never taught us this and threw these papers at us to do and i have no idea what any of this is

OpenStudy (anonymous):

cot(b) would be the reciprocal of that. so 2/\[\sqrt{5}\]. and then snce tanb = sin(b)/cos(b), sin(b)=\[\sqrt{5}\] and cos(b)=2.

OpenStudy (anonymous):

and also if csc(V) = (4*squareroote 7)/ 7 then what is sin (V) and cos (V) and tan (V)